Forces F1 (east) and F2 are simultaneously applied to a 3.0 Kg mass. When F2 is east, a = 5.0m/s2 east and when F2 is west a = 1.0m/s2 east. Determine the magnitude of F1 and F2.


1
Expert's answer
2021-06-16T16:38:16-0400
"F_2+F_1=ma_1=3(5)=15\\F_1-F_2=ma_2=3(1)=3\\
F_1=0.5(15+3)=9\ N\\F_2=0.5(15-3)=6\ N\\"